# A High Efficient and Scalable Obstacle-Avoiding VLSI Global Routing Flow

### v2

@@ 4.2 Rule Enhancements
(1) L-Shape Optimization: Although the edge updating performs well for the cases in Figure 2, there are still some cases that are difficult with only the basic edge updating. For example, In Figure 3(a), 3 steiner nodes from FLUTE are shown. Edge updating needs to update $[n_{0},n_{1}]$, and the solution is Figure 3(a) as the left corner of $b0$ is nearer to $[n_{0},n_{1}]$. However, if $n_{0}$ connects $n_{2}$ by the edges on the right of Figure 3(b), after $[n_{0},n_{1}]$ and a segment of $[n_{1},n_{2}]$ are removed, the optimal solution can be obtained. This can be realized by setting the reference line as $[n_{0},n_{2}]$. Besides, if we change the L-shape direction as Figure 3(c), another optimal solution can also be obtained. Inspired by these findings, given an edge $[n_{i},n_{j}]$ needs to be updated, if the degree of $n_{i}$ or $n_{j}$ is 2 (which means there exists an L-shape), the reference line of edge updating is the diagonal of this L-shape. Besides, two types of L-shapes are both tried to perform the edge updating, and the solution with a smaller wirelength is kept.
@@
(2) Reference Line Sloping: Besides for L-shapes, there are also other cases that can be explored. For example, Figure 3(d) shows the solution of the proposed edge updating for $[n_{0},n_{1}]$, but a better solution is shown in Figure 3(e). This can be obtained by setting the reference line with a slope to the left. Therefore, when updating $[n_{i},n_{j}]$, all its orthogonal half-lines (at most 4) starting from $n_{i}$ (or $n_{j}$) are considered to build the reference lines between the source node $n_{j}$ (or $n_{i}$) and a node on the half-lines, respectively. $k_{l}$ evenly distributed reference lines within the bounding box are tried for each half-line, where $k_{l}$ is set by users. An example is shown in Figure 3(f). After updating edge $[n_{0},n_{1}]$ by multiple reference lines, the solution with the smallest wirelength is kept.
@@
(3) Obstacle Merging: Another approach for rule enhancement is to merge multiple nearby obstacles, and the bounding box of these merged obstacles is regarded as a new obstacle when applying edge updating. In Figure 3(g), the original edge updating incurs large cost due to lots of bends, and the sloped reference line on the left can also incur large cost. In contrast, the solution in Figure 3(h) becomes more optimized where $b_{0}$ and $b_{1}$ are merged. Note that sometimes merging obstacles can cause overlaps, such as $b_{1}$ in Figure 3(i). Considering the runtime, if the edge updating is blocked by the overlapped obstacles, the edge extends on the boundaries of the overlapped obstacles, through the shortest path until able to perform the edge extension without the overlaps. Finally, similar to the reference line sloping approach, when updating each edge $[n_{0},n_{1}]$, $k_{m}$ merging strategies are tried, where $k_{m}$ is set by users. More specifically, if there are $n^{\prime}$ obstacles that cross $[n_{0},n_{1}]$, the merging strategies of merging each $1$, $n^{\prime}/k_{m}$, $2n^{\prime}/k_{m}$, $k_{m}\times n^{\prime}/k_{m}$ obstacles are tried, and the solution with the smallest wirelength is kept.
@@ 4.3 The Rule-Based OARSMT Algorithm
1:$S_{Fe}\leftarrow$ Run FLUTE to get the solution of edges
2:$S_{e}\leftarrow$ Build the rectilinear tree based on $S_{Fe}$
@@
15:     for Each merging strategy solution $S_{mb}$ of $S_{candb}$ do
16:         for Each possible $l_{r}$ and source node $n_{s}$ of $e$ do
17:              $U_{e}\leftarrow$ Edge_Update($n_{s},l_{r},S_{mb}$)
18:              if $WL(U_{e})<WL(U_{emin})$, $U_{emin}\leftarrow U_{e}$
19:         end for
20:     end for
21:     $S_{e}\leftarrow S_{e}\setminus e$
22:     $S_{e}\leftarrow S_{e}\bigcup U_{emin}$
23:end for
24:Remove redundant edges, merge overlapped edges in $S_{e}$
25:Return $S_{e}$
@@
(2) Steiner Nodes Legalization: The second step is to legalize the steiner nodes inside the obstacles (Lines 3-6) as Figure 2(c). For a set of invalid steiner nodes $S_{n}$ within an obstacle $b$, the intersection points between $b$’s boundaries and edges connecting nodes in $S_{n}$ are found. Then, the shortest edges placed around the obstacle boundaries that are able to connect all such intersections are calculated and added to edge set $S_{e}$. Finally the edge segments inside $b$ are removed from $S_{e}$. An example is also shown as the green parts in Figure 4.
(3) Iterative Edge Updating: The critical step is edge updating, which is iteratively performed (Lines 7-23). Each time an edge $e$ with violations is selected, the set of obstacles $S_{be}$ that block $e$ is obtained (Line 8). Then, a bounding box $B_{b}$ is heuristically introduced to greatly reduce the complexity (Line 12). The bounding box is a rectangle that just surrounds $e$ and all obstacles in $S_{be}$, and the candidate obstacles $S_{candb}$ are selected as the obstacles overlap $B_{b}$ (Line 13). For example, in Figure 4, when updating edge $[n_{0},n_{1}]$, $B_{b}$ is built based on $[n_{0},n_{1}]$, $b_{4}$ and $b_{5}$. In this case, $S_{candb}=\{b_{4},b_{5}\}$. In contrast, when updating edge $[p_{1},n_{3}]$, although $B_{b}$ is not built based on $b_{2}$, but $b_{2}$ is included in $S_{candb}$ since $B_{b}$ overlap $b_{2}$. Next, a record of the minimum wirelength connection is initiated at Line 14, and the edge updating by all combinations of approaches introduced in Sections 4.1 and 4.2 is performed to $e$ (Lines 15-16), with the consideration of the obstacles in $S_{candb}$. After this, the merged obstacles $S_{mb}$, a reference line $l_{r}$ and a source node $n_{s}$ is selected, and the edge updating is performed (Line 17). Once a solution with better wirelength is found, the record $U_{emin}$ is updated (Line 18). The best solution is selected to replace edge $e$ (Lines 21-22). Note that all updated new edges will be further checked after edge updating of $e$, as they might also cross obstacles, so finally all edges are legal.
(4) Post Processing: Finally, each edge connecting a steiner node with degree 1 is removed, and the overlapped edges are merged (Line 24). This can be finished quickly by checking the edges connected to each steiner node or pin. Additionally, we perform cycle detection on the generated tree. Even if a non-tree structure is produced, it can be quickly repaired using classical algorithms. In our experiments, the probability of generating a non-tree structure is $\sim$1%.
@@ 4.4 Data Structure Improvement and Time Complexity
The critical and the most time-consuming part is the iterative edge updating (Lines 7-23 in Algorithm 1). During this part, frequently checking for if a line crosses any obstacles (and which obstacles) is needed, such as by Line 8 of Algorithm 1, especially when bending an edge during the edge updating in Section 4.1. This can be realized by using an R-tree to store obstacles and querying it to check if a segment or rectangle overlaps with any obstacles. However, the R-tree structures are relatively complex and general that also allow the insertions and overlaps of the stored rectangles (obstacles). As obstacles are typically fixed and can be represented by non-overlap rectangles, the R-tree data structure can be replaced by simpler data structure for higher efficiency. Therefore, we propose a specific data structure to avoid redundant functions and it is illustrated in Figure 5. For example, Figure 5 shows a layout with obstacles indicated as gray rectangles. The Hanan grid can then be generated as the dotted lines. Note that the grid is not only based on the pins and the boundaries of obstacles, but also based on the centers of obstacles such as $k_{b3}$ of $b_{3}$. For each row and column of the grid, a range list is maintained with multiple sorted pairs of numbers, separated by semicolons. Each pair is a range of an obstacle’s width or height, respectively. We name each pair as range pair. For example, for Row 1 ($y=1$), there are two obstacles $b_{0}$ and $b_{1}$, which are located at $0\leq x\leq 3$ and $4\leq x\leq 5$, respectively. In this case, two range pairs are stored as $\{0,3;4,5\}$. Besides, to track which obstacle each range pair corresponds to, additional lists containing the obstacle pointer for each range pair are maintained, which is not shown in this example. The range lists are built in $O(n)$ runtime for one time, by checking the locations and ranges of all obstacles and pushing the ranges into the corresponding list.
@@
After the range lists are built, checking which obstacles a given line crosses can be quickly finished. The procedure can be illustrated by checking the line $e$ in Figure 5. $e$ is horizontally located at Row 2, so the x-coordinates (1 and 6) of $e$’s two nodes $n_{0}$ and $n_{1}$ are checked in the range list $\{0,3;4,5;7,8\}$ to obtain the list indices where these two x-coordinates are. $n_{0}$ locates between $x=0$ and $x=3$ (between indices 0 and 1 of $\{0,3;4,5;7,8\}$), while $n_{1}$ locates between $x=5$ and $x=7$ (between indices 3 and 4). The obstacles of which the range pairs’ indices are between the indices of $n_{0}$’s and $n_{1}$’s locations are categorized as crossed the line $e$. In this example, the obstacles cross $e$ are $b_{0}$ and $b_{1}$. The obstacles can be obtained by the maintained obstacle pointers associated with the range pairs. The checking only needs $O(log(n))$, since each list is sorted and contains at most all $n$ obstacles’ ranges.
Furthermore, if all obstacles that a bounding box (such as $B_{b}$ in Figure 5) overlap need to be obtained, a short boundary of the bounding box ($[c_{0},c_{1}]$) can be selected, and each grid line orthogonal to the short boundary inside $B_{b}$ is then selected, which is shown as the 4 horizontal lines in $B_{b}$ (including 2 long boundaries and the line through the center of $b_{3}$). Finally, these selected lines are iteratively checked to get all overlapped obstacles. The runtime is $O(len_{minb}\times log(n))$, where $len_{minb}$ is the length of the short boundary of the bounding box.
Based on this, the overall time complexity is analyzed as follows. The main time-consuming part is edge updating. Since an edge at most crosses $n$ obstacles, the bending of the edge performs at most $n$ times. Each time an edge is bent, the crossed obstacles need to be checked and the first one to the source node is selected, so it needs $O(nlog(n))$ runtime in total for each edge updating. When $k_{l}$ for reference line sloping and $k_{m}$ for obstacle merging are globally fixed in the tool for any layouts as in our experiments, it does not affect the time complexity. In Algorithm 1, as there are $m$ pins, the order of the edges in $S_{e}$ from FLUTE is $O(m)$. This infers that Lines 7-23 basically perform $O(m)$ times. As Lines 8, 12, and 17 needs $O(log(n))$, $O(len_{minb}log(n))$, and $O(nlog(n))$ runtime, the overall time complexity is $O(m\times(log(n)+len_{minb}log(n)+n\times log(n)))$. In practice, usually the maximum size of the support layout for a placer is fixed, so $len_{minb}$ is a constant. In this case, the time complexity is $O(m\times n\times log(n))$. This is less than the state-of-the-art previous work [11]. Besides, in most cases, an edge crosses only a small portion of obstacles, and the $n$ in the time complexity is much smaller than the number of all obstacles. This further leads to fast speed in practice.

### v3

@@ 4.2 Rule Enhancements
The enhanced rules are derived from Algorithm 1 Edge_Update($n_{s},n_{t},l_{r},S_{b}$), with modifications to the parameters $n_{s}$, $n_{t}$, $l_{r}$, and $S_{b}$.
@@
ER1 L-Shape Optimization: Although edge updating performs well for the cases in Figure 2, there are still cases where basic edge updating alone is insufficient. For example, in Figure 3(a), an L-shape connection generated based on FLUTE is shown. Here, $n_{0}$ and $n_{2}$ are two tree nodes connected by an L-shaped tree edge, while $n_{1}$ represents a corner node for this connection. Edge updating needs to update $n_{0},n_{1}]$, and the solution is Figure 3(a) as the left corner of $b_{0}$ is closer to $n_{0},n_{1}]$. However, if $n_{0}$ connects $n_{2}$ by the edges on the right of Figure 3(b), after $n_{0},n_{1}]$ and a segment of $n_{1},n_{2}]$ are removed, the optimal solution can be obtained. This can be realized by setting the reference line as $n_{0},n_{2}]$. Moreover, changing the L-shape direction (where $n_{3}$ is also a corner node) as shown in Figure 3(c) might also yield another optimal solution, even when there is a new obstacle $b_{3}$. Inspired by these findings, given an edge $n_{i},n_{j}]$ needs to be updated, if the degree of $n_{i}$ or $n_{j}$ is 2 and the other incident edge is orthogonal to $n_{i},n_{j}]$ (which means there exists an L-shape), the reference line of edge updating is the diagonal of this L-shape. Besides, two types of L-shapes are both tried to perform the edge updating, and the solution with a smaller wirelength is kept.
@@
ER2 Reference Line Sloping: Besides for L-shapes, there are also other cases that can be explored. For example, Figure 3(d) shows the solution of the proposed edge updating for $n_{0},n_{1}]$, but a better solution is shown in Figure 3(e). This can be obtained by setting the reference line with a slope to the left. In this case, given a general edge $n_{i},n_{j}]$ (such as $n_{0},n_{1}]$ in Figure 3(f)) that needs to be updated, the orthogonal edges to $n_{i},n_{j}]$ from node $n_{i}$ or $n_{j}$ are checked. Note that if no orthogonal edge exists in a certain direction of $n_{i}$ (or $n_{j}$), the sloping reference line from $n_{j}$ (or $n_{i}$) in that direction is not considered, as this may introduce additional edges and lead to higher cost. Such as two black edges from $n_{0}$ to its left and right, and one edge from $n_{1}$ to its left in Figure 3(f). The ray segments covering the checked edges from $n_{i}$ or $n_{j}$ are obtained, such as $n_{0},left)$, $n_{0},right)$, and $n_{1},left)$. For each obtained ray segment, $k_{l}$ nodes in total are used for reference line building, such as $n_{0}^{\prime}$ to $n_{k-1}^{\prime}$. These $k_{l}$ nodes are named hook nodes. Next, multiple reference lines $l_{r}$ for edge updating are built from $n_{i}$ (or $n_{j}$) to the hook nodes on the obtained ray segments from $n_{j}$ (or $n_{i}$), respectively. Such as reference lines $n_{1},n_{0}^{\prime}]$, $n_{1},n_{1}^{\prime}]$, etc. Note that the spacing between each nearest pair of hook nodes on the same ray segment is the same. Besides, all hook nodes are within the bounding box of $n_{i},n_{j}]$ (the minimal rectangular region that encloses all obstacles overlapping with $n_{i},n_{j}]$, such as the gray rectangle in Figure 3(f)). Finally, each reference line $l_{r}$ built from $n_{i}$ (or $n_{j}$) to a hook node is separately tried for Edge_Update($n_{s}$$n_{t}$$l_{r}$$S_{b}$) with $n_{s}\leftarrow n_{i}$ (or $n_{j}$) and $n_{t}\leftarrow n_{j}$ (or $n_{i}$), respectively. The solution with the smallest local wirelength is kept.
@@
ER3 Obstacle Merging: Another approach for rule enhancement is to merge multiple nearby obstacles, and the bounding box of these merged obstacles is regarded as a new obstacle when applying edge updating. In Figure 3(g), the original edge updating incurs large cost due to lots of bends, and the sloped reference line on the left can also incur large cost. In contrast, the solution in Figure 3(h) becomes more optimized where $b_{0}$ and $b_{1}$ are merged. Note that sometimes merging obstacles can cause overlaps, such as $b_{1}$ in Figure 3(i). For efficiency, if the edge updating is blocked by the overlapped obstacles, the edge extends on the boundaries of the overlapped obstacles, through the shortest path until able to perform the edge extension without the overlaps. Finally, totally $k_{m}+1$ ways of merging are tried for each edge $n_{s},n_{t}]$ to be updated as follows, where $k_{m}$ is user defined parameter. Assume there are $n^{\prime}$ successive obstacles $b_{0},b_{1},...b_{n^{\prime}-1}$ crossing edge $n_{s},n_{t}]$, then each group of $n_{m}$ successive obstacles can be temporarily merged. For example, if $n_{m}=3$, the obstacles in each group, such as $\{b_{0},b_{1},b_{2}\}$, $\{b_{3},b_{4},b_{5}\}$, $\{b_{6},b_{7},b_{8}\}$, etc., are separately merged as one obstacle. The value $n_{m}$ is tried multiple times as $1$, $n^{\prime}/k_{m}$, $2n^{\prime}/k_{m}$, $k_{m}\times n^{\prime}/k_{m}$. For each $n_{m}$ value tried, a temporary new obstacle set $S_{b}$ is obtained, and the edge updating for $n_{s},n_{t}]$ by Edge_Update($n_{s}$$n_{t}$$l_{r}$$S_{b}$) is performed. The solution with the smallest local wirelength among $k_{m}+1$ times trying is kept.
When updating an edge, the 3 enhanced rules are combined to be tried and the solution with the smallest cost (wirelength) is kept. This combination will be introduced in Algorithm 2 later.
@@ 4.3 The Rule-Based OARSMT Algorithm
[…]
@@
15:  if $e$ is on an L-shape pattern $L$ then //ER1
16:   $l_{r}\leftarrow$ the diagonal of $L$
17:   $n_{s},n_{t}\leftarrow$ 2 end nodes of $e$
18:   $U_{e}\leftarrow$ Edge_Update($n_{s},n_{t},l_{r},S_{candb}$)
19:   $L^{\prime}\leftarrow$ the L-shape by flipping $L$ to another direction
20:   $e^{\prime}\leftarrow$ the edge on $L^{\prime}$, with a same end node as $e$
21:   Calculate $U_{e^{\prime}}$ by performing Lines 8, 12-13, 16-18 by replacing $e$ with $e^{\prime}$
22:   if $WL(U_{e})<WL(U^{\prime}_{e})$, $U_{emin}\leftarrow U_{e}$ else ($U_{emin}\leftarrow U^{\prime}_{e}$ and $e\leftarrow e^{\prime}$)
23:  else
24:   for Each output $S_{mb}$ from ER3 when merging $S_{candb}$ do //ER3
25:     for Each possible $l_{r}$ and the source-target node pair ($n_{s},n_{t}$) of $e$ by ER2 do //ER2
26:      $U_{e}\leftarrow$ Edge_Update($n_{s},n_{t},l_{r},S_{mb}$)
27:      if $WL(U_{e})<WL(U_{emin})$, $U_{emin}\leftarrow U_{e}$
28:     end for
29:   end for
30:  end if
31:  $S_{e}\leftarrow S_{e}\setminus e$
32:  $S_{e}\leftarrow S_{e}\bigcup U_{emin}$
33:end for
34:Remove redundant edges, merge overlapped edges in $S_{e}$
35:Return $S_{e}$
@@
(2) Nodes Legalization: The second step is to legalize the nodes (Steiner nodes and L-shape corner nodes) inside the obstacles (Lines 3-6) as Figure 2(e). For a set of invalid nodes $S_{n}$ within an obstacle $b$, the intersection points between $b$’s boundaries and edges connecting nodes in $S_{n}$ are found. Then, the shortest edges placed around the obstacle boundaries that are able to connect all such intersections are calculated and added to edge set $S_{e}$. Finally, the edge segments inside $b$ are removed from $S_{e}$. An example is also shown as the green parts in Figure 4.
(3) Iterative Edge Updating: The critical step is edge updating, which is iteratively performed (Lines 7-33). Each time an edge $e$ with violations is selected, the set of obstacles $S_{be}$ that block $e$ is obtained (Line 8). Then, a bounding box $B_{b}$ is heuristically introduced to greatly reduce the complexity (Line 12). The bounding box is a rectangle that just surrounds $e$ and all obstacles in $S_{be}$, and the candidate obstacles $S_{candb}$ are selected as the obstacles overlap $B_{b}$ (Line 13). For example, in Figure 4, when updating edge $n_{0},n_{1}]$, $B_{b}$ is built based on $n_{0},n_{1}]$, $b_{4}$ and $b_{5}$. In this case, $S_{candb}=\{b_{4},b_{5}\}$. In contrast, when updating edge $p_{1},n_{3}]$, although $B_{b}$ is not built based on $b_{2}$, but $b_{2}$ is included in $S_{candb}$ since $B_{b}$ overlap $b_{2}$. Next, a record of the minimum wirelength connection is initiated at Line 14. The enhanced rule ER1 is first checked at Line 15. If $e$ is on an L-shape pattern, the reference line $l_{r}$, the source node $n_{s}$, and the target node $n_{t}$ are extracted as Lines 16-17. The edge updating from Algorithm 1 is called to obtain the updated edges $U_{e}$. Then, according to ER1, the edge of another direction of L-shape is tried to be updated (Lines 19-21), and the better solution is kept (Line 22). Otherwise, if $e$ is not on an L-shape, the edge updating is conducted (Line 26) by trying all combinations of $n_{s}$, $n_{t}$, $l_{r}$, and $S_{mb}$ from ER2 and ER3 (Lines 24-25)
Once a solution with better wirelength is found, the record $U_{emin}$ is updated (Line 27). The best solution is selected to replace edge $e$ (Lines 31-32). Note that all updated new edges will be further checked after edge updating of $e$, as they might also cross obstacles, so finally all edges are legal.
(4) Post Processing: Finally, each edge connecting a node (excluding pins) with degree 1 is removed, and the overlapped edges are merged (Line 34). This can be finished quickly by checking the edges connected to each node (a Steiner node, an L-shape corner node, or a pin). Additionally, we perform cycle detection on the generated tree. Even if a non-tree structure is produced, it can be quickly repaired using classical algorithms. In our experiments, the probability of generating a non-tree structure is $\sim$1%.
@@ 4.4 Data Structure Improvement and Time Complexity
The critical and the most time-consuming part is the iterative edge updating (Lines 7-23 in Algorithm 2). During this part, frequently checking for if a line crosses any obstacles (and which obstacles) is needed, such as by Line 8 of Algorithm 2, especially when bending an edge during the edge updating in Section 4.1. This can be realized by using an R-tree to store obstacles and querying it to check if a segment or rectangle overlaps with any obstacles. However, the R-tree structures are relatively complex and general that also allow the insertions and overlaps of the stored rectangles (obstacles). As obstacles are typically fixed and can be represented by non-overlap rectangles, the R-tree data structure can be replaced by a simpler data structure for higher efficiency. Therefore, we propose a specific data structure to avoid redundant functions and it is illustrated in Figure 5. For example, Figure 5 shows a layout with obstacles indicated as gray rectangles. The Hanan grid can then be generated as the dotted lines. Note that the grid is not only based on the pins and the boundaries of obstacles, but also based on the centers of obstacles such as $k_{b3}$ of $b_{3}$. For each row and column of the grid, a range list is maintained with multiple sorted pairs of numbers, separated by semicolons. Each pair is a range of an obstacle’s width or height, respectively. We name each pair as range pair. For example, for Row 1 ($y=1$), there are two obstacles $b_{0}$ and $b_{1}$, which are located at $0\leq x\leq 3$ and $4\leq x\leq 5$, respectively. In this case, two range pairs are stored as $\{0,3;4,5\}$. Besides, to track which obstacle each range pair corresponds to, additional lists containing the obstacle pointer for each range pair are maintained, which is not shown in this example. The range lists are built in $O(n)$ runtime for one time, by checking the locations and ranges of all obstacles and pushing the ranges into the corresponding list.
@@
After the range lists are built, checking which obstacles a given line crosses can be quickly finished. The procedure can be illustrated by checking the line $e$ in Figure 5. $e$ is horizontally located at Row 2, so the x-coordinates (1 and 6) of $e$’s two nodes $n_{0}$ and $n_{1}$ are checked in the range list $\{0,3;4,5;7,8\}$ to obtain the list indices where these two x-coordinates are. $n_{0}$ locates between $x=0$ and $x=3$ (between indices 0 and 1 of $\{0,3;4,5;7,8\}$), while $n_{1}$ locates between $x=5$ and $x=7$ (between indices 3 and 4). The obstacles of which the range pairs’ indices are between the indices of $n_{0}$’s and $n_{1}$’s locations are categorized as crossed the line $e$. In this example, the obstacles cross $e$ are $b_{0}$ and $b_{1}$. The obstacles can be obtained by the maintained obstacle pointers associated with the range pairs. This check only takes $O(log(n))$ time, since each list is sorted and contains at most all $n$ obstacle ranges.
Furthermore, if all obstacles that a bounding box (such as $B_{b}$ in Figure 5) overlap need to be obtained, a short boundary of the bounding box ($c_{0},c_{1}]$) can be selected, and each grid line orthogonal to the short boundary inside $B_{b}$ is then selected, which is shown as the 4 horizontal lines in $B_{b}$ (including 2 long boundaries and the line through the center of $b_{3}$). Finally, these selected lines are iteratively checked to get all overlapped obstacles. The runtime is $O(len_{minb}\times log(n))$, where $len_{minb}$ is the length of the short boundary of the bounding box.
Based on this, the overall time complexity is analyzed as follows. The main time-consuming part is edge updating. Since an edge at most crosses $n$ obstacles, the bending of the edge performs at most $n$ times. After each time of bending, we identify the 1st obstacle intersecting the edge’s current extending path (Line 5 of Algorithm 1), thus requiring $O(nlog(n))$ runtime in total for each edge updating. When $k_{l}$ for reference line sloping and $k_{m}$ for obstacle merging are globally fixed in the tool for any layouts as in our experiments, it does not affect the time complexity. In Algorithm 2, as there are $m$ pins, the order of the edges in $S_{e}$ from FLUTE is $O(m)$. This implies that Lines 7–23 are essentially executed $O(m)$ times. As lines 8, 12, and 17 needs $O(log(n))$, $O(len_{minb}log(n))$, and $O(nlog(n))$ runtime, the overall time complexity is $O(m\times(log(n)+len_{minb}log(n)+n\times log(n)))$. In practice, usually the maximum size of the support layout for a placer is fixed, so $len_{minb}$ is a constant. In this case, the time complexity is $O(m\times n\times log(n))$. This is less than the state-of-the-art previous work [11]. Besides, in most cases, an edge crosses only a small portion of obstacles, and the $n$ in the time complexity is much smaller than the number of all obstacles. This further leads to fast speed in practice.